calculate the pressure exerted by 10.2 two of NH3 in a 3. 0 DM cube vessel at 25 degree Celsius by using ideal gas equation

Grade:11

1 Answers

Khimraj
5 years ago
mole of NH3 = 10.2/17 = 0.6
volume = 3L
temperature = 298K
P = nRT/V = 0.6*0.0821*298/3 = 4.89 atm
Hope it clears.............................................